OpenStudy (anonymous):

Solve for y in the following system of equations. -2x - 2y = -12 2x - 2y = 28 can some one please help me

OpenStudy (anonymous):

I just need to know how to work it out not the answer.

OpenStudy (anonymous):

first you want to find for y...so because there is a -2 and a positive 2x x is out of the equation then you want to add -2y and -2y which is -4y then you will add -12 and 28 which is 16....last you divide 16 by -4 to find out y

OpenStudy (anonymous):

then you substitute that answer in for y and do the same to find x...if you need more help let me know

OpenStudy (anonymous):

ok thank you message me if u want bored in clASS

OpenStudy (texaschic101):

-2x - 2y = - 12 2x - 2y = 28 -------------add 0 - 4y = 16 -4y = 16 y = - 16/4 y = 4 and if you need to find x, just sub 4 in for y in either of the original equations, and solve for x. Once you know what x and y are, you can check your answer by subbing in the known values into either of the equations and if they come out equal, then you are correct.